In a certain school, the students engage in various sports in the following
proportions:
Cricket (C) : 60% of all students
Tennis (T) : 50% of all students
Both Cricket & Tennis : 30% of all students
If a student is selected at random, what is the probability that he will
(i) Play Cricket or Tennis
(ii) Play neither sports?
1
Expert's answer
2012-06-22T08:10:35-0400
Cricket only: Co = C - B = 60% - 30% = 30%,
Tennis only: To = T - B = 50% -
30% = 20%.
So, 100% - Co - To - B = 100% - 30% - 20% - 30% = 20% of students
play
neither sports.
So, 100% - 20% = 80% play Cricket or
Tennis.

Answer: (i) 80%, (ii) 20%.